7. If f(x) = (2x2 + 3x - 7), find f(x).
Answer #1

solution a - 3 --) [= 요 dar (-)*) 6 (+37 - 4 a 2717 -3) da S = 6 (37-) - (2 (27) +3 . - વિન3) ટ